# Deformation quantization and perverse sheaves

### v1

@@ Acknowledgements
We would like to thank Claude Sabbah, Pierre Schapira and Nick Sheridan for useful conversations.
@@ 1.2. Contactification
It is not in general possible to find a contactification with $\rho\colon\Lambda\rightarrow L$ an analytic isomorphism everywhere, see [DK11, Example 2.5.4].
@@
The existence of such a contactification is shown in [DK11, Proposition 2.5.3]. Let us now show that any two such contactifications are uniquely isomorphic. By [Gie64] we may choose a triangulation of $S$ with $L\subset S$ as a simplicial subcomplex. Therefore, there is a neighborhood $U$ of $L$ which deformation retracts to $L$.
@@
Possibly refining the cover, we choose holomorphic functions $g_{i}$ on $U_{i}$ such that $dg_{i}=\theta_{i}$. Then $g_{i}-g_{j}-\varphi_{ij}$ define locally constant functions on $U_{ij}$ which determine a cohomology class in $\mathrm{H}^{1}(U;\mathbf{C})$. The restriction of these functions to $L$ is $(g_{i}-f_{i})-(g_{j}-f_{j})$. Note that $d(g_{i}-f_{i})|_{L_{\mathrm{reg}}}=0$, i.e. $g_{i}-f_{i}$ is locally constant on $L_{\mathrm{reg}}\cap U_{i}$ and hence, by density, on $L\cap U_{i}$ (see [DK11, Sublemma 2.5.2]). Thus, under the restriction map $\mathrm{H}^{1}(U;\mathbf{C})\rightarrow\mathrm{H}^{1}(L;\mathbf{C})$ the class $[g_{i}-g_{j}-\varphi_{ij}]$ restricts to zero. Since $U$ deformation retracts to $L$, this implies that there are locally constant functions $h_{i}$ on $U_{i}$ such that
@@ 1.3. Orientation torsors
which we call the Maslov quadratic form. For $k=\mathbf{R}$ its signature is the Maslov index $\tau(L_{1},L_{2},L_{3})$ of the triple of Lagrangians as defined in [KS94, Chapter A.3].
@@ 1.4. Charts on symplectic manifolds
For the statement on the level of symplectic manifolds, see [Wei71, Theorem 7.1]. By Example 1.7 $j^{1}0\subset\mathrm{J}^{1}L$ is a contactification of the zero section $L\subset\mathrm{T}^{*}L$ and by 1.12 contactifications of Lagrangians are unique. This implies the statement about the local form of the contactification. ∎
@@
It is not in general possible to arrange $f$ to be zero on $L\cap M$ when this intersection is non-reduced, see [Joy15, Example 2.13].
@@
Consider the map $\Xi\colon L\rightarrow L\times M$ from 1.29. By construction the critical loci of $f$ on $L$ and $h$ on $L\times M$ are both $L\cap M$. So, by [Joy15, Proposition 2.25] there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{L\cap M}$ of the embedding $\Xi\colon L\rightarrow L\times M$. It has the following alternative description. For any $x\in L\cap M$ we have a sequence $\mathrm{T}_{M,x},\mathrm{T}_{\pi_{1},x},\mathrm{T}_{\pi_{2},x},\mathrm{T}_{M,x}$ of Lagrangian subspaces of $\mathrm{T}_{S,x}$, where the subsequent subspaces are transverse. Therefore, we have the Maslov quadratic form $q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})$ on $\mathrm{T}_{M}|_{L\cap M}$.
@@ 2.1. Monodromic $D$-modules
If $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$, then $[\mathcal{M}]$ is a regular holonomic $\mathcal{D}_{X}$-module with an endomorphism $\theta$. For brevity we call such a module a regular holonomic $\mathcal{D}_{X}[\theta]$-module. The following statement is standard.
@@
  • $\mathcal{M}$ is lisse along the orbits of the $\mathbf{C}^{\times}$-action.
@@ 2.2. Monodromic perverse sheaves
  • The stack of perverse sheaves on $X\times\mathbf{C}^{\times}$ lisse along $\mathbf{C}^{\times}$.
@@
    on $Y$ is lisse along $\mathbf{C}^{\times}$.
@@
We will also consider perverse sheaves of $\mathbf{C}[\![\hbar]\!]$-modules as in [DGS11, Section 5]. We say a perverse sheaf $\mathcal{L}\in\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$ is torsion-free if the map $\hbar\colon\mathcal{L}\rightarrow\mathcal{L}$ is a monomorphism in the abelian category $\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$.
@@
2.3. The Riemann-Hilbert functor
@@
$\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\to\mathrm{Perv}_{\mathrm{diff}}(X)$
@@
By [KS12, Proposition 7.1.7] a constructible complex $\mathcal{F}$ is cohomologically complete, i.e.
@@
is an isomorphism is the content of [PSY14, Theorem 7.8] in the case $X=\mathrm{pt}$ and the same proof works for sheaves. ∎
We are now ready to define the functor $\widehat{\mathrm{RH}}^{-1}$. Given a perverse sheaf $\mathcal{F}$ equipped with an automorphism $T$ we can write $T=\exp(-2\pi iM)$ for a unique endomorphism $M\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G\subset\mathbf{C}$. Then we define $\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$ to be the pair $(\mathcal{F}(\!(\hbar)\!),D)$, where
@@
The functor $\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\rightarrow\mathrm{Perv}_{\mathrm{diff}}(X)$ is fully faithful. Moreover, if $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(U)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, then RH^1(/)superscriptnormal-^normal-$\mathcal{M}\cong\widehat{\mathrm{RH}}^{-1}(\mathcal{L}/\hbar)$. In particular, the essential image of $\widehat{\mathrm{RH}}^{-1}$ consists of differential perverse sheaves with regular singularities.
@@
Given a morphism $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}$ in $\mathrm{Perv}_{\mathrm{aut}}(X)$, a morphism $\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{1})\rightarrow\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{2})$ of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves is uniquely specified by its restriction $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ to $\mathcal{F}_{1}\hbar^{0}\subset\mathcal{F}_{1}(\!(\hbar)\!)$. Let $\mathcal{F}_{1}^{\lambda}\subset\mathcal{F}_{1}$ be a generalized eigenspace of $D$ with eigenvalue $m\in G$, where $\exp(-2\pi im)=\lambda$. The map $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ will therefore send $\mathcal{F}_{1}^{\lambda}$ to the $m$-generalized eigenspace of 2(())subscript2Planck-constant-over-$\mathcal{F}_{2}(\!(\hbar)\!)$ which is precisely $\mathcal{F}_{2}^{\lambda}\hbar^{0}\subset\mathcal{F}_{2}(\!(\hbar)\!)$. The resulting map $\mathcal{F}_{1}^{\lambda}\rightarrow\mathcal{F}_{2}^{\lambda}$ is compatible with the nilpotent part of $D$ precisely when it is compatible with the unipotent part of $T$. This proves fully faithfulness.
@@
is an isomorphism, so this identifies $\mathcal{L}\cong\mathcal{F}[\![\hbar]\!]$. Therefore, $\mathcal{E}\cong\mathcal{F}(\!(\hbar)\!)=\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$. Under this isomorphism the action of $D$ on $\mathcal{E}$ coincides with the action of $D$ coming from the monodromy operator $T$ on $\mathcal{F}$. ∎
@@ 3.1. Microdifferential operators
In this section we recall some algebras of microdifferential operators that we will use. We refer to [Kas03, Kas86] for more details. Recall the punctured cotangent bundle $\mathring{\mathrm{T}}^{*}X=\mathrm{T}^{*}X\setminus\mathrm{T}^{*}_{X}X$ and the projectivized cotangent bundle $\mathrm{P}^{*}X=\mathring{\mathrm{T}}^{*}X/\mathbf{C}^{\times}$ with $\gamma\colon\mathring{\mathrm{T}}^{*}X\rightarrow\mathrm{P}^{*}X$ the natural projection. We denote the projections $\mathrm{T}^{*}X\rightarrow X$ and $\mathrm{P}^{*}X\rightarrow X$ by the same letter $\pi$.
Consider the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators introduced in [SKK73]. It has a filtration
@@
Sometimes one considers the sheaf $\mathcal{E}_{X}$ of microdifferential operators; the sheaf ^$\widehat{\mathcal{E}}_{X}$ is obtained from $\mathcal{E}_{X}$ by completion with respect to the order. The theory of regular holonomic $\mathcal{E}_{X}$- and ^$\widehat{\mathcal{E}}_{X}$-modules coincide (see e.g. [KK83, Theorem 6.3]), so the difference is inessential for our purposes.

### v2

@@ Acknowledgements
We would like to thank Claude Sabbah, Pierre Schapira and Nick Sheridan for useful conversations. S.G. was partially supported by NSF grant DMS-2202363.
@@ 1.2. Contactification
It is not in general possible to find a contactification with $\rho\colon\Lambda\rightarrow L$ an analytic isomorphism everywhere, see [DAgnoloKashiwara, Example 2.5.4].
@@
The existence of such a contactification is shown in [DAgnoloKashiwara, Proposition 2.5.3]. Let us now show that any two such contactifications are uniquely isomorphic. By [Giesecke] we may choose a triangulation of $S$ with $L\subset S$ as a simplicial subcomplex. Therefore, there is a neighborhood $U$ of $L$ which deformation retracts to $L$.
@@
Possibly refining the cover, we choose holomorphic functions $g_{i}$ on $U_{i}$ such that $dg_{i}=\theta_{i}$. Then $g_{i}-g_{j}-\varphi_{ij}$ define locally constant functions on $U_{ij}$ which determine a cohomology class in $\mathrm{H}^{1}(U;\mathbf{C})$. The restriction of these functions to $L$ is $(g_{i}-f_{i})-(g_{j}-f_{j})$. Note that $d(g_{i}-f_{i})|_{L_{\mathrm{reg}}}=0$, i.e. $g_{i}-f_{i}$ is locally constant on $L_{\mathrm{reg}}\cap U_{i}$ and hence, by density, on $L\cap U_{i}$ (see [DAgnoloKashiwara, Sublemma 2.5.2]). Thus, under the restriction map $\mathrm{H}^{1}(U;\mathbf{C})\rightarrow\mathrm{H}^{1}(L;\mathbf{C})$ the class $[g_{i}-g_{j}-\varphi_{ij}]$ restricts to zero. Since $U$ deformation retracts to $L$, this implies that there are locally constant functions $h_{i}$ on $U_{i}$ such that
@@ 1.3. Orientation torsors
which we call the Maslov quadratic form. For $k=\mathbf{R}$ its signature is the Maslov index $\tau(L_{1},L_{2},L_{3})$ of the triple of Lagrangians as defined in [KashiwaraSchapiraSheaves, Chapter A.3].
@@ 1.4. Charts on symplectic manifolds
For the statement on the level of symplectic manifolds, see [WeinsteinLagrangian, Theorem 7.1]. By Example 1.7 $j^{1}0\subset\mathrm{J}^{1}L$ is a contactification of the zero section $L\subset\mathrm{T}^{*}L$ and by 1.12 contactifications of Lagrangians are unique. This implies the statement about the local form of the contactification. ∎
@@
It is not in general possible to arrange $f$ to be zero on $L\cap M$ when this intersection is non-reduced, see [JoycedCrit, Example 2.13].
@@
Consider the map $\Xi\colon L\rightarrow L\times M$ from 1.29. By construction the critical loci of $f$ on $L$ and $h$ on $L\times M$ are both $L\cap M$. So, by [JoycedCrit, Proposition 2.25] there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{L\cap M}$ of the embedding $\Xi\colon L\rightarrow L\times M$. It has the following alternative description. For any $x\in L\cap M$ we have a sequence $\mathrm{T}_{M,x},\mathrm{T}_{\pi_{1},x},\mathrm{T}_{\pi_{2},x},\mathrm{T}_{M,x}$ of Lagrangian subspaces of $\mathrm{T}_{S,x}$, where the subsequent subspaces are transverse. Therefore, we have the Maslov quadratic form $q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})$ on $\mathrm{T}_{M}|_{L\cap M}$.
@@ 2.1. Monodromic $D$-modules
If $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$, then $[\mathcal{M}]$ is a regular holonomic $\mathcal{D}_{X}$-module with an endomorphism $\theta$. For brevity we call such a module a regular holonomic $\mathcal{D}_{X}[\theta]$-module. The following statement is standard (see e.g. [Brylinski, Proposition 7.12]).
@@
  • $\mathcal{M}$ is locally constant along the orbits of the $\mathbf{C}^{\times}$-action.
@@ 2.2. Monodromic perverse sheaves
  • The stack of perverse sheaves on $X\times\mathbf{C}^{\times}$, which are locally constant along the $\mathbf{C}^{\times}$ orbits.
@@
    on $Y$ is locally constant along the $\mathbf{C}^{\times}$ orbits.
@@
We will also consider perverse sheaves of $\mathbf{C}[\![\hbar]\!]$-modules as in [DAgnoloGuillermouSchapira, Section 5]. We say a perverse sheaf $\mathcal{L}\in\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$ is torsion-free if the map $\hbar\colon\mathcal{L}\rightarrow\mathcal{L}$ is a monomorphism in the abelian category $\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$.
@@
We will need the following Yoneda lemma for monodromic perverse sheaves. Consider a $\mathbf{C}^{\times}$-torsor $\gamma\colon\tilde{Y}\rightarrow Y$ and let
$\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\subset\mathrm{Perv}(\tilde{Y})$
be the full substack of perverse sheaves, which are locally constant along the fibers of $\gamma$, and
$\mathrm{LocSys}(\tilde{Y})\subset\mathrm{Perv}(\tilde{Y})$
the full substack of local systems. There is a natural action of the symmetric monoidal stack $\mathrm{LocSys}(Y)$ on $\mathrm{LocSys}(\tilde{Y})$ (via $\gamma^{*}$) and $\mathrm{Perv}(Y)$. For $\mathcal{F}\in\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ the derived pushforward $\gamma_{*}\mathcal{F}$ has cohomology only perverse degrees $-1$ and 00.
Lemma 2.7.
Consider the functor
$\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\longrightarrow\mathrm{Fun}_{\mathrm{LocSys}(Y)}(\gamma_{*}\mathrm{LocSys}(\tilde{Y}),\mathrm{Perv}(Y))$
of stacks on $Y$ given by sending $\mathcal{F}\in\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ to the $\mathrm{LocSys}(Y)$-linear functor $V\in\mathrm{LocSys}(\tilde{Y})\mapsto{}^{p}\mathcal{H}^{-1}(\gamma_{*}(V\otimes\mathcal{F}))$. It induces an isomorphism after applying $\mathbb{R}\mathcal{H}om(-,-)$: i.e. for every $\mathcal{F}_{1},\mathcal{F}_{2}\in\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ the natural morphism
$\mathbb{R}\mathcal{H}om(\mathcal{F}_{1},\mathcal{F}_{2})\longrightarrow\mathbb{R}\mathcal{H}om({}^{p}\mathcal{H}^{-1}(\gamma_{*}((-)\otimes\mathcal{F}_{1})),{}^{p}\mathcal{H}^{-1}(\gamma_{*}((-)\otimes\mathcal{F}_{2})))$
is an isomorphism of complexes of sheaves on $Y$.
Proof.
The question is local on $Y$, so we may assume that $\tilde{Y}=Y\times\mathbf{C}^{\times}\rightarrow Y$ is a trivial $\mathbf{C}^{\times}$-torsor and $Y$ is contractible. By 2.1 both $\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\cong\mathrm{Perv}_{\mathrm{aut}}(Y)$ and $\gamma_{*}\mathrm{LocSys}(\tilde{Y})$ split into blocks according to the generalized eigenvalue of the monodromy operator. This allows us to reduce the statement to a similar statement for the substack of perverse sheaves $\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)\subset\mathrm{Perv}_{\mathrm{aut}}(Y)$ consisting of pairs $(\mathcal{F},T)\in\mathrm{Perv}_{\mathrm{aut}}(Y)$, where the automorphism $T\colon\mathcal{F}\rightarrow\mathcal{F}$ is unipotent, i.e. $(T-\mathrm{id})^{N}=0\colon\mathcal{F}\rightarrow\mathcal{F}$ for some $N$. Moreover, since $Y$ is assumed to be contractible, $\gamma_{*}\mathrm{LocSys}(\tilde{Y})$ is given by the stack of $\mathbf{C}[T^{\pm 1}]$-modules. For a local system $V$ of $\mathbf{C}[T^{\pm 1}]$-modules and $\mathcal{F}\in\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)$ we have
${}^{p}\mathcal{H}^{-1}(\gamma_{*}(V\otimes\mathcal{F}))\cong(V\otimes\mathcal{F})^{T},$
where on the right we consider $T$-invariants. This sheaf vanishes if the $T$-action on $V$ is not unipotent, so we will further restrict to such modules.
So, for $\mathcal{F}_{1},\mathcal{F}_{2}\in\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)$ we have to show that the natural morphism
$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]}(\mathcal{F}_{1},\mathcal{F}_{2})\longrightarrow\mathbb{R}\mathcal{H}om_{\mathbf{C}}(((-)\otimes\mathcal{F}_{1})^{T},((-)\otimes\mathcal{F}_{2})^{T})$
is an isomorphism. Fix a natural number $N$, so that $(T-1)^{N}$ annihilates both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. In this case the left-hand side becomes isomorphic to
$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}(\mathcal{F}_{1},\mathcal{F}_{2})$
and the right-hand side becomes isomorphic to
$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}((\mathbf{C}[T^{\pm 1}]/(T-1)^{N}\otimes\mathcal{F}_{1})^{T},(\mathbf{C}[T^{\pm 1}]/(T-1)^{N}\otimes\mathcal{F}_{2})^{T}),$
which is again isomorphic to $\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}(\mathcal{F}_{1},\mathcal{F}_{2})$ by an explicit computation of the invariants. ∎
2.3. The Riemann–Hilbert functor
@@
$\mathrm{RH}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\to\mathrm{Perv}_{\mathrm{diff}}(X)$
@@
By [KashiwaraSchapira, Proposition 7.1.7] a constructible complex $\mathcal{F}$ is cohomologically complete, i.e.
@@
is an isomorphism is the content of [PSY, Theorem 7.8] in the case $X=\mathrm{pt}$ and the same proof works for sheaves. ∎
We are now ready to define the functor $\mathrm{RH}^{-1}$. Given a perverse sheaf $\mathcal{F}$ equipped with an automorphism $T$ we can write $T=\exp(-2\pi iM)$ for a unique endomorphism $M\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G\subset\mathbf{C}$. Then we define $\mathrm{RH}^{-1}(\mathcal{F})$ to be the pair $(\mathcal{F}(\!(\hbar)\!),D)$, where
@@
The functor $\mathrm{RH}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\rightarrow\mathrm{Perv}_{\mathrm{diff}}(X)$ is fully faithful. Moreover, if $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(U)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, then $\mathcal{M}\cong\mathrm{RH}^{-1}(\mathcal{L}/\hbar)$. In particular, the essential image of $\mathrm{RH}^{-1}$ consists of differential perverse sheaves with regular singularities.
@@
Given a morphism $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}$ in $\mathrm{Perv}_{\mathrm{aut}}(X)$, a morphism $\mathrm{RH}^{-1}(\mathcal{F}_{1})\rightarrow\mathrm{RH}^{-1}(\mathcal{F}_{2})$ of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves is uniquely specified by its restriction $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ to $\mathcal{F}_{1}\hbar^{0}\subset\mathcal{F}_{1}(\!(\hbar)\!)$. Let $\mathcal{F}_{1}^{\lambda}\subset\mathcal{F}_{1}$ be a generalized eigenspace of $D$ with eigenvalue $m\in G$, where $\exp(-2\pi im)=\lambda$. The map $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ will therefore send $\mathcal{F}_{1}^{\lambda}$ to the $m$-generalized eigenspace of 2(())subscript2Planck-constant-over-$\mathcal{F}_{2}(\!(\hbar)\!)$ which is precisely $\mathcal{F}_{2}^{\lambda}\hbar^{0}\subset\mathcal{F}_{2}(\!(\hbar)\!)$. The resulting map $\mathcal{F}_{1}^{\lambda}\rightarrow\mathcal{F}_{2}^{\lambda}$ is compatible with the nilpotent part of $D$ precisely when it is compatible with the unipotent part of $T$. This proves fully faithfulness.
@@
is an isomorphism, so this identifies $\mathcal{L}\cong\mathcal{F}[\![\hbar]\!]$. Therefore, $\mathcal{E}\cong\mathcal{F}(\!(\hbar)\!)=\mathrm{RH}^{-1}(\mathcal{F})$. Under this isomorphism the action of $D$ on $\mathcal{E}$ coincides with the action of $D$ coming from the monodromy operator $T$ on $\mathcal{F}$. ∎
@@ 3.1. Microdifferential operators
In this section we recall some algebras of microdifferential operators that we will use. We refer to [KashiwaraDmodules, KashiwaraIntroduction] for more details. Recall the punctured cotangent bundle $\mathring{\mathrm{T}}^{*}X=\mathrm{T}^{*}X\setminus\mathrm{T}^{*}_{X}X$ and the projectivized cotangent bundle $\mathrm{P}^{*}X=\mathring{\mathrm{T}}^{*}X/\mathbf{C}^{\times}$ with $\gamma\colon\mathring{\mathrm{T}}^{*}X\rightarrow\mathrm{P}^{*}X$ the natural projection. We denote the projections $\mathrm{T}^{*}X\rightarrow X$ and $\mathrm{P}^{*}X\rightarrow X$ by the same letter $\pi$.
Consider the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators introduced in [SatoKawaiKashiwara]. It has a filtration
@@
In Section 8 we will also encounter the sheaf $\mathcal{E}_{X}$ of non-formal microdifferential operators, so that ^$\widehat{\mathcal{E}}_{X}$ is obtained by completion from $\mathcal{E}_{X}$.